# RevCD - Reversed Conditional Diffusion for Generalized Zero-Shot Learning

### v1

@@ 5 Conclusion
In this paper, we introduce a reversed Conditional Diffusion model (RevCD) and evaluate its performance against VAEs and GANs for zero-shot learning. This foundational research explores the largely untapped potential of diffusion-based models to generate unseen samples. Our RevCD model generates samples for classes that serve as prototypes for high-accuracy classification. By leveraging visual conditioning, our approach allows precise control over the generation process, outperforming other generative methods in settings with unseen classes. Experimental results demonstrate the advantages of using a diffusion model as a generative backbone, especially regarding its robustness to limited semantic information. We believe our findings can stimulate further exploration of diffusion models in generalized zero-shot learning (GZSL). Moreover, expanding cross-dataset evaluations in future zero-shot learning research could lead to the development of more resilient models.

### v2

@@ 5 Conclusion
In this paper, we introduce a reversed Conditional Diffusion model (RevCD) and evaluate its performance against state-of-the-art methods for generalized zero-shot learning. Additionally, by comparing with VAEs and GANs, our research explores the largely untapped potential of diffusion-based models to generate unseen samples. Our RevCD model estimates the semantic density for classes, which is then used to generate prototypes of both seen and unseen instances for high-accuracy classification. By leveraging visual conditioning, our approach enables precise control over the generation process and improved posterior approximations, outperforming other generative methods in settings with seen classes. Experimental results demonstrate the advantages of using a diffusion model as a generative backbone, especially regarding its robustness to diverse semantic information. We believe our findings can stimulate further exploration of diffusion models in generalized zero-shot learning (GZSL). Moreover, expanding cross-dataset evaluations in future zero-shot learning research could lead to the development of more resilient models.